A total of 915 tickets were sold for a game for a total of $1,130. If adult tickets sold for $2.00 and children's tickets sold for $1.00, how many of each kind of ticket were sold?
adult ticket=
Children ticket=

Found 2 solutions by mananth, ewatrrr:
Answer by mananth(16949)

let adults be x
children be y
..
child ticket = 1
adult ticket =2
...
x+y =915
2x+y=1130
multiply first equation by -1
-x-y=-915
add it to second equation
x=215 adults
y=700 children
...
m.ananth@hotmail.ca

Answer by ewatrrr(24785)


Hi,
Let x represent the number of adult tickets sold. then (915-x) is the number of student tickets sold
write as we read***
$2*x + $1(915-x) = $1,130
solve for x
x = 1130-915
x = 215
215 adult tickets and 700 children's tickets were sold
checking our answer
$430 + $700 = $1,130